In Desmos, type in y= square root of x ; y= x2 in the graph


the domain of square root of x [0, infinity) and the range is [0, infinity)

the domain of x2 is ( - infinity, infinity) and the range is [0, infinity)


What do you notice about the domain/ range of square root of x and x2?


1
Expert's answer
2021-03-11T12:28:43-0500

given that :

1). "y=\sqrt{x}"


domain = "[0,\infty)" and range = "[0,\infty)"


2). "y=x^2"


domain = "(-\infty,+\infty)" and range = "[0,\infty)"


both graph have different domain but have same range, this is because as in first graph "\sqrt{x}" is present which is taken only the positive values of x for obtaining the real values range of graph. Thus, the domain of this graph is "[0,\infty)" and gives values as range is "[0,\infty)" .


But in the second graph x2 is present and due to square nature it can take all real values either positive or negative as domain but its range is "[0,\infty)", as all the negative values of the x turn to positive value due to squaring so it doesn't gives negative values.